# Resource-constrained Fairness

### v4

@@ 4 What influences the Cost of Fairness?
Base rate disparity
First, we alter the base rate disparity. We see in Figures 2(a) and 2(b), that a reduction in the base rate disparity leads to a lower cost of fairness. This trend is more pronounced when enforcing DP compared to EO, but it is evident under both fairness metrics. Menon and Williamson (2018) also find that the trade-off between accuracy and fairness is determined by the strength of the correlation between the sensitive attribute and the target.
Noise
We alter the performance of the model by introducing random noise (with varying degrees from 0 to 0.5) to the feature values of the whole dataset. The base rates of both groups remain the same, but the model struggles more with correctly classifying individuals, resulting in a lower AUC score. We see that as model performance decreases, the average cost of fairness also goes down, both for DP and EO (Figures 2(c)-2(d)). This global decrease shrinks the gap between the performances of each group, reducing the overall cost of fairness. This phenomenon might also explain why the average cost of fairness is so high for the CelebA dataset in Table 2. The model is very good in distinguishing negatives and positives from each other (as measured by the AUC), so the average cost will be high. [^6] Following this reasoning, we see that if we would fit a perfect model (so the model predicts the target label perfectly for all instances) to a dataset with some level of base rate disparity, then the maximum cost would be reached at a point between the base rates and be exactly equal to the level of base rate disparity in that dataset. In the case of CelebA, the model is not perfect but the maximum cost is also relatively close to the level of the base rate disparity, and is reached at a point between the base rates.
Subgroup noise
We can also add noise only to the members of the protected group. This will lower the performance of the model for individuals of the protected group, but not for individuals of the privileged group. As shown in Figures 2(e) and 2(f), this modification increases the cost of fairness for both DP and EO, with a larger effect for EO. Chen et al. (2018) also find that when there is a difference in noise level, and available covariates are not equally predictive of the outcome in both groups, fairness cannot be satisfied without sacrificing accuracy. Dutta et al. (2020) confirm that if there is not enough separability information for one group compared to the other, being fair will reduce the accuracy.
Size of the protected group
Lastly, the size of the protected group substantially influences the cost of fairness. When the protected group is smaller, fewer adjustments are necessary to achieve fairness, thereby reducing the cost. This relationship is clearly illustrated in Figures 2(g) and 2(h), where reducing the size of the protected group, by randomly excluding part of this group, consistently lowers the cost of fairness. This is an expected finding, however, we have not seen it discussed previously.
4.1 The impact of selection rate
@@
Figures 3(a)-3(f) demonstrate the loss in precision for various selection rates when using a fair allocation compared to the default allocation of the machine learning model. We see that this cost varies substantially depending on the available resources. For the Adult (Figure 3(a)), Compas (Figure 3(b)), CelebA (Figure 3(e)) and Fitzpatrick 17K (Figure 3(f)) datasets, the cost is highest for low selection rates, while for the Dutch (Figure 3(c)) dataset, the cost is the highest for medium selection rates, and for the Law (Figure 3(d)) dataset, the cost is the highest for high selection rates. The influence of resource level on the cost of fairness has not been discussed before, yet it is a critical factor. Depending on the resource level, the cost can either be negligible or very high. The loss in recall can be seen in Figure 5. We see that the patterns are in line with the loss in precision. We can also partially agree with Rodolfa et al. (2021) who posited that for very constrained top-$k$ settings, so when the selection rate is very low, the cost of fairness can be negligible. However, we only find this to be true when the base rate is a lot higher and the performance of the model is good enough. When the base rate is low, the cost of fairness will usually be high for low-resource settings.
(a) Selection rate = 10 %
(b) Selection rate = 25 %
(c) Selection rate = 50 %
(d) Selection rate = 75 %
Figure 4: Allocation results for different resource levels (Adult Income dataset). Left point on the x-axis represents the resources being maximally allocated to the privileged group, while the right point represents the resources being maximally allocated to the protected group.
We can get more insights into how these trade-offs behave for each resource level by analyzing the different scenarios. For the Adult dataset, this is shown in Figures 4(a)-4(d). These figures demonstrate how different allocations lead to different levels of precision for one fixed resource level. The two extreme points of the curve represent all the resources being awarded to either the privileged group (red dot on the left) or the protected group (purple dot on the right).  [^7] However, also for these extreme points, when the total level of resources is higher than the size of one of the groups, some of the resources will still be awarded to the other group (as can for example be seen in Figure 4(c)). We see that the optimum is not reached by awarding all the resources to one of the groups, but somewhere in the middle (darkblue dot). We note that the precision of the unconstrained (‘unfair’) model, will be very close to the optimal allocation, and that the allocations required by the fairness metrics (both DP and EO) will lead to a lower precision for every resource level (Figures 4(a)-4(d)). However, this difference is a lot larger for low resource levels (Figures 4(a)-4(b)). In Figure 4(c), the default allocation of the machine learning model is very unfair, but using the fair allocation (both DP and EO) results in only a little change in precision. This demonstrates that a very unfair initial model allocation does not guarantee a high cost of fairness. In Figure 4(d), the default allocation of the machine learning model is already approximately fair. Hence, the cost of fairness is also low.
@@ 4.2 The Impact of the Fairness Metric
We can see in Table 2, and in Figures 34 and 5 that the cost of fairness is usually lower when enforcing equal opportunity than when enforcing demographic parity. This is because the number of instances that needs to be ‘swapped’ is typically lower when enforcing equal opportunity. Liu et al. (2018) also find that the selection rate enforced by equal opportunity is likely to be much closer to the optimal than the selection rate enforced by demographic parity, while Hardt et al. (2016) show that enforcing equal opportunity leads to less loss in profits than enforcing demographic parity.
The setting where EO would results in a higher number of swapped instances, is when an equal true positive rate results in stricter requirements than an equal positive rate. A case where this could happen is when one of the groups has significantly lower data quality. DP suffers less under this scenario, as here we do not care if the predictions are accurate or not. However, to enforce EO, the algorithm must do equally bad on all groups, leading to a worse overall performance. Indeed, as shown in Figures 2(e) and 2(f), the cost of enforcing EO increases significantly more rapidly compared to the cost of enforcing DP when one subgroup becomes noisier.
4.3 Bounding the Cost of Fairness
We now formalize these costs and consider the cost of the change induced by enforcing fairness at a particular selection rate.
As we are modeling the cost of inducing fairness under rate constraints, we consider the changes in the labeling induced by swaps, where the label assigned to one instance changes from positive to negative, and the label assigned to another instance changes from negative to positive. A sequence of swaps must always preserve the selection rate, and any pair of labellings with the same selection rate, can be transformed from one to another by a sequence of swaps.
We write $p$ for the proportion of instances in the dataset swapped from positive to negative (or equivalently the proportion swapped from negative to positive). The overall cost of inducing fairness can be written as:
$\text{cost}=p\times\text{average cost of a swap}$ (7)
Note that $p$ is somewhere between 0 and 0.5, and for standard linear measures (e.g. accuracy, recall, sensitivity, and False Positive and Negative Rates) the second term lies somewhere between $-c^{-1}$, and $c^{-1}$. Here,
$c=\frac{\text{\#instances used in the measure}}{\text{\#instances in dataset}}$ (8)
i.e. $c=1$ for accuracy, and, writing $b$ for the base rate, $c=b$ for Recall and False Positive Rate, and $c=(1-b)$ for Specificity and False Negative Rate. As such, $c^{-1}p$, and half the average cost of a swap are upper bounds on the cost of enforcing fairness.
To state this less formally, the cost of fairness, in terms of change in accuracy, is at most the proportion of swaps ($p$); while the cost in terms of change of recall is at most the proportion of swaps divided by the base rate.
We can create successive bounds based on this identity.
Given a fixed global selection rate $r$, $p$ is bounded above by $r$ (all positive instances are swapped); and $(1-r)$ (all negative instances are swapped). Therefore, the cost of fairness is bounded above by $rc^{-1}$, and $(1-r)c^{-1}$. As such, if $r$ or $(1-r)$ is much smaller than $c$, there is little cost in enforcing fairness.
This matches our empirical finding that the cost is bell-shaped with respect to the global selection-rate. The highest costs occur when the selection rates are closest to the selection rates of an unconstrained classifier. This means that many of the existing works on the cost of fairness (Menon and Williamson, 2018; Friedler et al., 2019; Hort et al., 2023) overestimate the cost of fairness for resource-constrained scenarios. For scenarios where the selection rate substantially exceeds the base-rate, this is because for any informative classifier, the majority of candidates that might be positive would be selected early on, regardless of which group they belong to, and the remaining pool of candidates are likely to have negative labels regardless of which group they belong to.
Similarly, given $g$, the proportion of the dataset occupied by the smallest group, we know $p<g$, and the cost of inducing fairness is bounded above by $gc^{-1}$. [^8] Relabelling the smallest group is not always sufficient to enforce fairness, but it holds for the fairness metrics of demographic parity and equal opportunity that we study in this paper. Again, if $g$ is much smaller than $c$, there is little cost in inducing fairness. This result is supported empirically by our experiments (see Figures 2(g) and 2(h)).
Finally, looking at the second term (the average cost of relabelling), we find this second factor explains much of what we see. For example, increasing global noise decreases performance for both groups, at broadly similar proportions. This overall decrease shrinks the gap between the per-group performance, reducing the average cost of relabelling points. Similarly, increasing classification noise for the group with greatest classification error, or increasing the difference in base rates, increases the performance gap between groups and consequentially the cost of relabelling.
Bounding loss in precision
These bounds are much weaker for non-linear measures such as precision, if they exist. For precision, assuming a fixed global selection rate $r$, we have $c=r$. Consequentially, the low selection-rate bound $rc^{-1}=1$ is particularly meaningless. However, other bounds remain informative when there is a higher selection rate – particularly, the bounds $(1-r)c^{-1}$ and $gc^{-1}$, indicate that we should see similar patterns in the costs with respect to precision for higher selection rates. Empirically, we find this — precision exhibits similar trends to the linear measures of recall and accuracy. Reflecting the much weaker bound for low selection rates, it sometimes exhibits strong changes at low resource levels, which are not present for accuracy and recall. This can be seen in Figure 3(d).
4.4 Varying Global Selection Rates
Strategic adjustments in selection rate may be necessary depending on the outcomes of a cost-benefit analysis. Until now, we assumed the resource level to be fixed, but practitioners may find it feasible to increase or decrease this level, particularly if the resource levels were initially determined under an unfair model allocation. For instance, in healthcare, additional investments could be directed towards expanding screening facilities to counterbalance the decrease in true positives, effectively maintaining the detection rates prior to implementing fairness measures. Similarly, in educational settings such as student admission, institutions might adjust the size of admitted cohorts, either decreasing it to preserve the average quality, or increasing it to have the same number of graduates. Donahue and Kleinberg (2020) also discuss how increasing the level of available resources is a critical goal where advocacy and political action can play a key role.
Following section 4.3, it is reasonable to ask if bounds exist for levelling up, and if there are bounded increases in selection rate for a disadvantaged group guaranteed to remove unfairness (while keeping the selection rate of the advantaged group as-is). In brief, the answer is yes for demographic parity; for equal opportunity you may need to relabel all the data in a disadvantaged group; and levelling up is not always possible for equal precision. These results are so different from section 4.3 because the analysis of the previous section was advantaged by the stability of measures – we altered per group selection rates and hoped that global measures remained stable. In this new case, we are trying to alter per-group measures and the stability of measures is now a disadvantage.
For demographic parity, there are no stability concerns. Let $g$ be the proportion of the dataset corresponding to the disadvantaged group, and $r$ the global classifier selection rate. Then an upper bound for the selection rate excluding members of the disadvantaged group is $r/(1-g)$, and at most we require an additional $gr/(1-g)$ proportion of points to be positively labelled to obtain parity.
For equal opportunity, only instances assigned a positive ground-truth label alter the recall, and it is possible that a badly-trained classifier selects all negatively labeled instances first, and only then positively labelled points. If we consider a classifier like this, and a group with only one positively labelled point (in the ground truth), all points within the group must be selected to get any non-zero recall rate for that group. As such, $g$ is a bound.
For equal precision, we consider the same example, and note that while you must select all datapoints to obtain non-zero precision, the precision will be the base rate of the group, and it may be insufficient for equality with respect to the other group.
Another strategic approach could involve refining the precision for the protected groups by collecting more data or enhancing data quality, albeit at an initial cost. This investment might reduce the overall fairness cost, yielding a more profitable model over time. However, high quality data may simply not be available for all groups or infeasible to collect for a variety of legal, political, and practical reasons, meaning it cannot be assumed that collecting more or “better" data will necessarily solve inequalities in resource distribution (Perez, 2019; Pot et al., 2019; Wachter et al., 2020).
@@ 5 Conclusion
This work addresses the practical concern of enforcing fairness with a limited budget or, equivalently, as the selection rate/threshold varies. We demonstrated that standard methods for fairness do not enforce fairness for a varying threshold (Figure  2). We unified leveling up (Mittelstadt et al., 2023) and fairness under budgetary constraints, and have shown that the solutions for each approach coincide under constrained resources (Section 2.2). We have also shown how simple methods for setting per group thresholds can enforce fairness under these rate constraints (Section 3). Using this, we empirically measured the cost of fairness under rate constraints – in terms of decreased performance, or as an increase in global harm (Section 4); and derived theoretical bounds for these costs (Section 4.3). While the bounds are not intended to be tight, they are informative, and align with our empiric findings.
This allows us to investigate the actual cost of fairness, rather than changes in performance metrics that are heavily influenced by a change in the overall selection rate (Goethals et al., 2024). Compared to other fairness analyses, our analysis more closely aligns with real-world challenges, where resources are often fixed, and allows organizations to have reasonable expectations about what costs to expect and why. Furthermore, we demonstrated that the decision-making context matters, particularly the behavior of the classifier; the used fairness metric; and the level of available resources. This highlights the importance of evaluating the cost of fairness within the specific domain of use.
@@ Acknowledgments
Appendix
@@
Figure 5: Cost of fairness (loss in recall) for all datasets when enforcing demographic parity (DP) and equal opportunity (EO). We see that the cost of fairness for both metrics heavily depends on the available resource level and thus the used selection rate.

### v5

@@ 6.1 The Impact of Selection Rate
Figures 3(a)-3(f) demonstrate the loss in precision for various selection rates when using a fair allocation compared to the default allocation of the machine learning model. We see that this cost varies substantially depending on the available resources. For the Adult (Figure 3(a)), Compas (Figure 3(b)), CelebA (Figure 3(e)) and Fitzpatrick 17K (Figure 3(f)) datasets, the cost is highest for low selection rates, while for the Dutch (Figure 3(c)) dataset, the cost is the highest for medium selection rates, and for the Law (Figure 3(d)) dataset, the cost is the highest for high selection rates. The influence of resource level on the cost of fairness has not been discussed before, yet it is a critical factor. Depending on the resource level, the cost can either be negligible or very high. The results for recall and accuracy are in line with precision and can be found in Figures 4 and 5. We can also partially agree with Rodolfa et al. (2021) who posited that for very constrained top-$k$ settings, so when the selection rate is very low, the cost of fairness can be negligible. However, we only find this to be true when the base rate is a lot higher and the performance of the model is good enough. We get more insights into how these trade-offs behave for each resource level by analyzing the different scenarios in Section C.
@@ 6.2 The Impact of the Fairness Metric
Table 1 and Figures 36 show that the cost of fairness is generally lower when enforcing equal opportunity (EO) than demographic parity (DP) across all selection rates. This is because EO typically requires fewer instances to be “swapped." Liu et al. (2018) also find that the selection rate enforced by EO is closer to the optimal, while Hardt et al. (2016) show that enforcing EO leads to a lower loss in profits than enforcing DP. However, EO can be costlier when ensuring an equal true positive rate imposes stricter constraints than an equal positive rate, for example when one group has lower data quality. In such cases, DP is less affected since it disregards prediction accuracy, whereas EO requires equally poor recall across groups, degrading overall performance. As shown in Figures 2(e)- 2(f), the cost of enforcing EO rises significantly faster than the cost of enforcing DP when one subgroup becomes noisier.
@@ 7 Conclusion
This work addresses the practical concern of enforcing fairness with a limited budget or, equivalently, as the selection rate (or threshold) varies. We unified leveling up Mittelstadt et al. (2023) and fairness under budgetary constraints, and have shown that the solutions for each approach coincide under constrained resources (Section 3). We have also shown how simple methods for setting per group thresholds can enforce fairness under these rate constraints (Section 4). Using this, we empirically measured the cost of fairness under rate constraints – in terms of decreased performance, or as an increase in global harm; and derived theoretical bounds for these costs (Section 6). While the bounds are not intended to be tight, they are informative, and align with our empirical findings. This allows us to investigate the actual cost of fairness, rather than changes in performance metrics that are heavily influenced by a change in the overall selection rate Goethals and Calders (2024). Compared to other fairness works, our analysis more closely aligns with real-world challenges, where resources are often fixed, and allows organizations to have reasonable expectations about what costs to expect and why. Furthermore, we demonstrated that the decision-making context matters, particularly the behavior of the classifier; the used fairness metric; and the level of available resources.
@@ Acknowledgments
Appendix A Materials and Methods
A.1 Materials
We use several real-world datasets common in tabular fair machine learning Le Quy et al. [2022]. We also extend our findings to two datasets (CelebA Liu et al. [2015] and Fitzpatrick17k Groh et al. [2021]) from computer vision. The Adult Income dataset is taken from from the 1994 census data, and contains a prediction task identifying if an individual’s annual income surpasses \$50,000. The Compas dataset gathers demographic details and criminal records of defendants from Broward County to forecast the likelihood of reoffending within two years. The Dutch Census dataset from 2001 captures aggregated demographic data in the Netherlands, utilized to determine if an individual’s occupation falls into a high-level (prestigious) or low-level category. The Law Admission dataset contains data from a 1991 survey by the Law School Admission Council (LSAC) across 163 U.S. law schools, aimed at predicting a student’s success on the bar exam. The CelebA dataset comprises over 200,000 celebrity images, each annotated with 40 attribute labels ranging from hair color to emotions, along with 5 landmark locations. We use the attribute ‘wearing earrings’ as target label, as it is highly skewed towards the non Male class Wang et al. [2020] and thus has a significant base rate disparity. The Fitzpatrick17k dataset consists of approximately 17,000 dermatologist-curated skin lesion images, categorized across the Fitzpatrick skin type scale. We preprocess the data and binarize the labels following the approach of Zong et al. [2022]. We specify the protected and target attributes in Table 2. We also report the base rate disparity for each dataset, which is defined as the actual difference in the proportion of positive outcomes between groups in a dataset. If we consider a binary sensitive attribute $S$, with $ns$ representing the disadvantaged group and $s$ the advantaged group it can be mathematically expressed as:
$P(Y=1\mid S=ns)-P(Y=1\mid S=s)$
[…]
A.2 Methods
For the tabular datasets, we use eXtreme Gradient Boosting (XGBoost) Chen and Guestrin [2016]. We optimize the number of boosting rounds through 5-fold cross-validation on the training set with early stopping after 10 rounds if no improvement is seen. The optimized model is trained on the entire training set and evaluates the probability of positive class outcomes. For the image datasets, we employ a Resnet-50 (CelebA) and Resnet-18 (Fitzpatrick-17k) backbone He et al. [2016] pretrained on ImageNet Deng et al. [2009] for feature extraction.
A.2.1 Comparison with other bias mitigation methods
Figure 1 illustrates how demographic parity, one of the most common notions of fairness, varies with selection rate, when deploying a set of popular bias mitigation methods. We use Disparate Impact Remover (DIR) Feldman et al. [2015], Adversarial Debiasing (ADV) Zhang et al. [2018] and Meta Fair Classifier (MFC) Celis et al. [2019], and use the implementation of AIF360 Bellamy et al. [2018]. The points represent the labels of the bias mitigation methods at their default threshold. However, as the thresholds vary, fairness oscillates wildly, revealing that these methods may break when considering the whole range of possible selection rates. For our experiments, we use a kind of threshold optimizer, which enforces fairness exactly at each possible selection rate.
A.3 Metrics
A.3.1 Performance metrics
$P(S(x_{i})>S(x_{z})\mid y_{i}=1,y_{z}=0)$
While precision at the top $R$ relates to the efficient use of limited resources, knowing this value also determines the values of both recall and accuracy at $R$ for a specific dataset and model and a fixed value of $R$ Rodolfa et al. [2021]. We measure the cost of fairness as the loss in precision when using the fair allocation instead of the optimal allocation. For the sake of completeness, we also report the cost as loss in recall and loss in accuracy in Table 1 and Figure 4. The formula to calculate recall at the top R:
Recall at R=Number of actual positives in the top RTotal number of actual $\text{Recall at $R$}=\frac{\text{Number of actual positives in the top $R$}}{\text{Total number of actual positives}}$
The formula to calculate accuracy at the top R:
Accuracy at R=Number of actual positives in the top R+Number of actual negatives not in the top RTotal $\text{Accuracy at $R$}=\frac{\text{Number of actual positives in the top $R$}+\text{Number of actual negatives not in the top $R$}}{\text{Total population}}$
We focus on these widely used performance metrics, but the results hold for any weighted sum of costs of Type I/II errors.
A.3.2 Fairness Metrics
We focus on two standard fairness measures used to assess disparities between groups.
Demographic parity (also known as statistical parity) requires that the rate of positive decisions is roughly equal for both the disadvantaged group and the advantaged group Dwork et al. [2012]:
$P(\hat{Y}=1|S=s)\approx P(\hat{Y}=1|S=ns)$
Equal opportunity requires the true positive rate to be approximately the same across groups Hardt et al. [2016], which enforces equal recall:
$P(\hat{Y}=1|S=s,Y=1)\approx P(\hat{Y}=1|S=ns,Y=1)$
Appendix B Results for other performance metrics
In this Section, we present the results for recall and accuracy. The general trends are in line with the results for precision.
@@ Appendix B Results for other performance metrics
Figure 4: Cost of fairness (loss in recall) for all datasets when enforcing demographic parity (DP) and equal opportunity (EO). The cost of fairness for both metrics heavily depends on the available resource level and thus the used selection rate.
(a) Adult
(b) Compas
(c) Dutch
(d) Law
(e) CelebA - Earrings
(f) Fitzpatrick17k
Figure 5: Cost of fairness (loss in accuracy) for all datasets when enforcing demographic parity (DP) and equal opportunity (EO). The cost of fairness for both metrics heavily depends on the available resource level and thus the used selection rate.
Table 3: The average cost of fairness (the loss in recall and accuracy) when enforcing DP and EO. The average is calculated over all the selection rates from $1\%$ to $100\%$.
Dataset Adult Compas Dutch Law CelebA Fitzpatrick17K
Avg. loss in recall (DP) 0.023 0.016 0.028 0.004 0.062 0.001
Avg. loss in recall (EO) 0.010 0.007 0.008 0.002 0.013 0.001
Avg. loss in accuracy (DP) 0.011 0.017 0.026 0.007 0.026 0.000
Avg. loss in accuracy (EO) 0.005 0.008 0.008 0.004 0.005 0.000
Appendix C Results for different resource levels
(a) Selection rate = 10 %
(b) Selection rate = 25 %
(c) Selection rate = 50 %
(d) Selection rate = 75 %
Figure 6: Allocation results for different resource levels (Adult Income dataset). Left point on the x-axis represents the resources being maximally allocated to the advantaged group, while the right point represents the resources being maximally allocated to the disadvantaged group.
We study the effect of the different resource level on the allocation of resources. For the Adult dataset, this is shown in Figures 6(a)-6(d). These figures demonstrate how different allocations lead to different levels of precision for one fixed resource level. The two extreme points of the curve represent all the resources being awarded to either the advantaged group (red dot on the left) or the disadvantaged group (purple dot on the right). [^8] However, also for these extreme points, when the total level of resources is higher than the size of one of the groups, some of the resources will still be awarded to the other group (as can for example be seen in Figure 6(c)). We see that the optimum is not reached by awarding all the resources to one of the groups, but somewhere in the middle (darkblue dot). We note that the precision of the unconstrained (‘unfair’) model, will be very close to the optimal allocation, and that the allocations required by the fairness metrics (both DP and EO) will lead to a lower precision for every resource level (Figures 6(a)-6(d)). However, this difference is a lot larger for low resource levels (Figures 6(a)-6(b)). In Figure 6(c), the default allocation of the machine learning model is very unfair, but using the fair allocation (both DP and EO) results in only a little change in precision. This demonstrates that a very unfair initial model allocation does not guarantee a high cost of fairness. In Figure 6(d), the default allocation of the machine learning model is already approximately fair. Hence, the cost of fairness is also low.